# ReALM: Reference Resolution As Language Modeling

### v1

@@ Abstract
Reference resolution is an important problem, one that is essential to understand and successfully handle context of different kinds. This context includes both previous turns and context that pertains to non-conversational entities, such as entities on the user’s screen or those running in the background. While LLMs have been shown to be extremely powerful for a variety of tasks, their use in reference resolution, particularly for non-conversational entities, remains underutilized. This paper demonstrates how LLMs can be used to create an extremely effective system to resolve references of various types, by showing how reference resolution can be converted into a language modeling problem, despite involving forms of entities like those on screen that are not traditionally conducive to being reduced to a text-only modality. We demonstrate large improvements over an existing system with similar functionality across different types of references, with our smallest model obtaining absolute gains of over 5% for on-screen references. We also benchmark against GPT-3.5 and GPT-4, with our smallest model achieving performance comparable to that of GPT-4, and our larger models substantially outperforming it.
@@ 1 Introduction
Here, it is immediately apparent that it would not be possible for the Agent to understand or complete the user’s query without the ability to use and comprehend context. It also stands to reason that there are multiple types of context that are necessary to handle user queries: conversational context and on-screen context being two prime examples.
Recent Large Language Models (LLMs) Stammbach et al. (2022); Touvron et al. (2023); Santhanam et al. (2022); Dettmers et al. (2023) have often enabled end-to-end experiences, perhaps even obviating the need of a traditional multi-stage pipeline that includes reference resolution. There are, however, still several real-world cases where a pipeline is valuable, perhaps even essential, and an end-to-end approach falls short. First, when a framework runs completely on-device (for example, for privacy and efficiency reasons) on a system such as a smartphone that has relatively limited computing power, due to the low-power nature of the system and latency constraints, using a single, large, end-to-end model is infeasible: using a single LLM for this task would usually require the use of a large model with long prompts for true end-to-end experiences Wei et al. (2022). Second, consider the case when the model has to integrate with APIs, has to consume information from components upstream, or has to provide information to be consumed downstream: while in these cases it is possible to have an end-to-end approach having the LLM write API calls Patil et al. (2023); Qin et al. (2023), this often requires a large language model and a complete overhaul of existing pipelines, which might be cumbersome or completely infeasible. Third, the use of a focused model would allow for an existing reference resolution module to be swapped with improved versions in a transparent way, while providing improved ability to hill-climb and improved interpretability, by virtue of the system being modular. Finally, for the task under consideration in this paper, reference resolution does not include solely conversational references, but also includes the ability to reference an on-screen and/or a background entity that is part of what the user currently perceives in their interaction with a device, but has not been a part of the conversational history that results from their direct interaction with the virtual agent in question. There thus continues to be utility in exploring "traditional" NLP tasks such as reference resolution, despite some of the larger language models being able to handle them implicitly. In this work, we thus advocate the use of (relatively) smaller language models, but fine-tuned for specifically and explicitly for the task of reference resolution.
@@
2 Related Work and Motivation
While traditional reference resolution systems have explored conversational and visual/deictic references in great depth Kottur et al. (2018); Schwartz et al. (2019); Kang et al. (2019), resolving on-screen references is a domain that has been relatively under-explored. However, as shown above, conversational agents on a mobile device need to understand references to the screen, and to support such experiences, to be truly natural. On screen references differ from visual and deictic references for several reasons: they tend to be more structured and highly textual, which enables the use of a lighter model to treat this as a text-only problem without a visual component; further, user queries around on-screen elements often tend to be more action-oriented rather than QA based; finally, they use synthetic screens rather than natural real-world images, which are much easier to parse, but whose distribution completely differs from that on which larger pre-trained image-based systems (such as CLIP Radford et al. (2021)) tend to be trained. Further, jointly being able to perform conversational and on-screen has been even less explored.
@@ 2 Related Work and Motivation
The most closely related work which we are aware of, and which we consequently use as our baseline, is that of Ates et al. (2023), an extension of Bhargava et al. (2023) which deals purely with on-screen references; however, it suffers from several drawbacks, which we address in this work. First, this approaches rely on a dedicated “Category module” to deal with type-based references. This module often requires manually on-boarding entities every time a new type is created (a common occurrence in voice assistants, as the supported functionality of assistant is expanded over time). In addition, such modules often treat each type as distinct, with the similarity of different types ignored. This, in turn, leaves on the table the potential positive transfer that could have happened between semantically related classes (such as “phone number” and “contact”) when data is added for one of those classes. This approach is thus difficult to scale to new entity types and use cases. Second, these systems rely on the use of hand-crafted rule-based textual overlap features, which require heavy feature engineering and tend not to be robust. In addition, these heuristics often do not account for semantic similarity, and are not able to encode real-world understanding or commonsense reasoning. Finally, these methods effectively classify how related each entity is to the query in question independently of all other entities, and neither take into consideration the whole screen nor the other entities.
@@ 3 Task
    Conversational Entities: These are entities relevant to the conversation. These entities might come from a previous turn for the user (for example, when the user says “Call Mom”, the contact for Mom would be the relevant entity in question), or from the virtual assistant (for example, when the agent provides a user a list of places or alarms to choose from).
@@
We pose the task of reference resolution as a multiple choice task for the LLM, where the intended output is a single option (or multiple options) from the entities shown on the user’s screen. In some cases, the answer could also be "None of these".
To evaluate this task, we allow the model to output the relevant entities in any order, i.e. if the Ground Truth is entities 8, 7, and 4, then we accept any permutation of these 3 correct entities while evaluating the performance of the model.
@@ 4 Datasets
Our datasets comprise data that was either synthetically created, or created with the help of annotators. Each data point contains the user query and a list of entities, along with the ground-truth entity (or set of entities) that are relevant to the corresponding user query. Each entity, in turn, contains information about its type and other properties such as the name and other textual details associated with the entity (the label and time of an alarm, for example). For data points where relevant on-screen context exists, this context is available in the form of the bounding box of the entity, and the list of objects surrounding it along with properties of these surrounding objects such as their types, textual contents and locations.
@@ 4.1 Conversational Data
In this case, data is collected for entities that are relevant to the user interaction with the agent. To do this, graders are shown screenshots with synthetic lists of entities provided, and asked to provide queries that unambiguously references an arbitrarily picked entity in the synthetic list provided. For example, graders might be provided a synthesized list of businesses or alarms and asked to refer to a particular entity within that list.
For example, the grader might be shown a list of businesses that are synthetically constructed, and they then refer to a specific one in the list provided. For example, they might say “Take me to the one that’s second from the bottom” or “Call the one on Main Street”.
@@ 4.2 Synthetic Data
There are two templates to generate the synthetic data. The first “base” template includes mentions, entities, and possible slot values if necessary. The second “language” template template imports the base template and adds different variations of queries that can be used for targeted cases with references defined in the base template.
The data generation script takes the base and language templates and generates the possible queries given in language template by substituting references with mention and slot values defined in base template. It loops through all supported entities. For the ones matching with the entity in the template, it connects mention and the entity, otherwise it only adds entity types with no reference to mention.
4.3 On-screen Data
Screen data were collected from various web pages where phone number, e-mail and/or physical address information exist. We had two phased annotation processes on screen data. First phase was for extracting queries based on the screens and second one was for identifying the entities and mention for the given query. In the first grading project, graders were given a screenshot (Figure 1(a)) with green and red boxes, the information in the green box and asked to classify the green boxed data into one of the entities such as phone number, email address, etc. Then, graders were asked to provide three unique queries for the green boxed data.
@@ 4.3 On-screen Data
Figure 1: Sample screenshots used in annotation
In second annotation project (Figure 1(b)), queries collected in first step were shown to graders one by one with corresponding screenshots without bounding boxes and all screen entities as a list. The graders were asked if the query contains mention to the one of the given visual entities, if the query sound natural. Also, they were asked to provide the entities from the list referred in the given query and tag the part of the query referring that entity.
@@ 5 Models
We compare our proposed model ReALM, described in detail in Section 5.3 below, with two baseline approaches: one based on a re-implementation of the reference resolver proposed in MARRS (Section 5.1), which is non-LLM based, and one based on ChatGPT (both GPT-3.5 and GPT-4; Section 5.2).
@@ 5.3 Our Approach
For each data point consisting of a user query and the corresponding entities, we convert it to a sentence-wise format that we can feed to an LLM for training. Examples of the input before and after processing are shown in Appendix Sections A and C, with examples of how we convert entities of different types into text shown in Appendix B. Note that the entities are shuffled before being sent to the model so that the model does not overfit to particular entity positions.
@@ 5.3.1 Conversational References
For the sake of this work, we assume conversational references to be of two types: type-based and descriptive. Type-based references are heavily reliant on using the user query in conjunction with the types of the entities to identify which entity (of a set of entities) are most relevant to the user query in question: for example, if the user says “play this”, we know that they are referring to an entity like a song or a movie, as opposed to a phone number or an address; “call him” likewise refers to the first among a set of phone numbers or contacts, as opposed to an alarm. Descriptive references, in contrast, tend to use an property of the entity to uniquely identify it: “The one in Times Square” for example might help uniquely refer to one among a set of addresses or business. Note that it is often the case that references might rely on both types and descriptions to unambiguously refer to a single object: consider the examples “play the one from Abbey Road” vs “directions to the one on Abbey Road”, both of which rely on both the entity type and description to identify a song in the first case and address in the second. In our proposed approach, we simply encode the type and various properties of the entity. We show our detailed encoding scheme in Appendix B.
@@ 6 Results
Table 3: Model Accuracy for Different Datasets. A prediction is correct is the model correctly predicts all relevant entities, and incorrect otherwise. Conv refers to the Conversational Dataset, Synth to the Synthetic one, Screen to the Onscreen one and Unseen to a conversational dataset pertaining to a held-out domain.
@@
Additionally, we also experiment with models of different sizes. We see that while performance improves in all datasets with an increase in model size, the difference is most pronounced for the onscreen datasets, which alludes to the task being more complex in nature.

### v2

@@ Abstract
Reference resolution is an important problem, one that is essential to understand and successfully handle contexts of different kinds. This context includes both previous turns and context that pertains to non-conversational entities, such as entities on the user’s screen or those running in the background. While LLMs have been shown to be extremely powerful for a variety of tasks, their use in reference resolution, particularly for non-conversational entities, remains underutilized. This paper demonstrates how LLMs can be used to create an effective system to resolve references of various types, by showing how reference resolution can be converted into a language modeling problem, despite involving forms of entities like those on screen that are not traditionally conducive to being reduced to a text-only modality. We demonstrate large improvements over an existing system with similar functionality across different types of references, with our smallest model obtaining absolute gains of over 5% for on-screen references. We also benchmark against GPT-3.5 and GPT-4, with our smallest model achieving performance comparable to that of GPT-4, and our larger models substantially outperforming it.
@@ 1 Introduction
Here, it is immediately apparent that it would not be possible for the Agent to understand or complete the user’s query without the ability to use and comprehend context. It also stands to reason that there are multiple types of context that are necessary to handle user queries: conversational context, on-screen context, and background entities.
Recent Large Language Models (LLMs) Stammbach et al. (2022); Touvron et al. (2023); Santhanam et al. (2022); Dettmers et al. (2023) have often enabled end-to-end experiences, perhaps even obviating the need of a traditional multi-stage pipeline that includes reference resolution Khatri et al. (2018). There are, however, still several real-world cases where a pipeline is valuable, perhaps even essential, and an end-to-end approach falls short. First, when a framework runs completely on-device (for example, for privacy and efficiency reasons) on a system such as a smartphone that has relatively limited computing power, due to the low-power nature of the system and latency constraints, using a single, large, end-to-end model is infeasible: using a single LLM for this task would usually require the use of a large model with long prompts for true end-to-end experiences Wei et al. (2022). Second, consider the case when the model has to integrate with APIs, has to consume information from components upstream, or has to provide information to be consumed downstream: while in these cases it is possible to have an end-to-end approach having the LLM write API calls Patil et al. (2023); Qin et al. (2023), this often requires a large language model and a complete overhaul of existing pipelines, which might be cumbersome or completely infeasible. Third, the use of a focused model would allow for an existing reference resolution module to be swapped with improved versions in a transparent way, while providing improved ability to hill-climb and improved interpretability, by virtue of the system being modular. Finally, for the task under consideration in this paper, reference resolution does not include solely conversational references, but also includes the ability to reference an on-screen and/or a background entity that is part of what the user currently perceives in their interaction with a device, but has not been a part of the conversational history that results from their direct interaction with the virtual agent in question. There thus continues to be utility in exploring "traditional" NLP tasks such as reference resolution, despite some of the larger language models being able to handle them implicitly. In this work, we thus advocate the use of (relatively) smaller language models, but fine-tuned for specifically and explicitly for the task of reference resolution.
@@
2 Realated Work and Motivation
While traditional reference resolution systems have explored conversational and visual/deictic references in great depth Kottur et al. (2018); Schwartz et al. (2019); Kang et al. (2019), resolving on-screen references is a domain that has been relatively under-explored. However, as shown above, conversational agents on a mobile device need to understand references to the screen, and to support such experiences, to be truly natural. On screen references differ from visual and deictic references for several reasons: they tend to be more structured and highly textual, which enables the use of a lighter model to treat this as a text-only problem without a visual component; further, user queries around on-screen elements often tend to be more action-oriented rather than QA based; finally, they use synthetic screens rather than natural real-world images, which are much easier to parse, but whose distribution completely differs from that on which larger pre-trained image-based systems (such as CLIP Radford et al. (2021)) tend to be trained. Further, jointly being able to perform conversational and on-screen reference resolution has been even less explored, with prior work often focusing on images and graphics Willemsen et al. (2023), or UI elements You et al. (2024).
@@ 2 Realated Work and Motivation
The most closely related work which we are aware of, and which we consequently use as our baseline, is that of Ates et al. (2023), an extension of Bhargava et al. (2023) which deals purely with on-screen references; however, it suffers from several drawbacks, which we address in this work. First, these approaches rely on a dedicated “Category module” to deal with type-based references. This module often requires manually on-boarding entities every time a new type is created (a common occurrence in voice assistants, as the supported functionality of the assistant is expanded over time). In addition, such modules often treat each type as distinct, with the similarity of different types ignored. This, in turn, leaves on the table the potential positive transfer that could have happened between semantically related classes (such as “phone number” and “contact”) when data is added for one of those classes. This approach is thus difficult to scale to new entity types and use cases. Second, these systems rely on the use of hand-crafted rule-based textual overlap features, which require heavy feature engineering and tend not to be robust. In addition, these heuristics often do not account for semantic similarity, and are not able to encode real-world understanding or commonsense reasoning. Finally, these methods effectively classify how related each entity is to the query in question independently of all other entities and later threshold them, whereas our current approach directly picks out the most relevant option (or options), while also allowing for no entities to be relevant. Our approach thus additionally has the advantage of removing the reliance on a set threshold, while also providing all the functionality supported in the previous approaches.
@@ 3 Task
    Conversational Entities: These are entities relevant to the conversation, which predominantly include those that come from a previous turn. For example, let’s say that the first turn of the user is “Call Mom.”, which is an unambiguous turn that uses a contact called Mom. Shortly after, if the user says “Text her”, the reference “her” needs to be resolved to the contact for “Mom” that was brought up in the previous turn; this contact is thus a conversational entity. Another example might involve an interaction in which the user requests for a list of places or alarms to choose from (or the agent presents one for a user turn such as “Show me pharmacies near me”); each item in this list then becomes a conversational entity for subsequent turns.
@@
We pose the task of reference resolution as a multiple choice task for the LLM, where the intended output is a single option (or multiple options) from the entities shown on the user’s screen. In some cases, the answer could also be "None of these", in which case the model needs to predict “0”.
To evaluate this task, we check if the predicted set of options matches the ground truth set; in other words, we allow the model to output the relevant entities in any order, i.e. if the Ground Truth is entities 8, 7, and 4, then we accept any permutation of these three correct entities while evaluating the performance of the model.
Note that as in Ates et al. (2023); Bhargava et al. (2023), we assume that entities along with their types come in from an upstream system (for example, through a mechanism involving entity pullers which are able to extract entities in a high recall manner or through a donation from a device app, as in Aas et al. (2023)).
@@ 4 Datasets
Our datasets comprise data that was either synthetically created, or created with the help of annotators. Each data point contains the user query and a list of entities, along with the ground-truth entity (or set of entities) that are relevant to the corresponding user query. Each entity, in turn, contains information about its type and other properties such as the name and other textual details associated with the entity (the label and time of an alarm, for example). For data points where relevant on-screen context exists, this context is available in the form of the bounding box of the entity, and the list of objects surrounding it along with properties of these surrounding objects such as their types, textual contents and locations. Note that our data collection follows that of Bhargava et al. (2023); Ates et al. (2023); we present an overview here and direct the interested reader to the aforementioned papers for a more detailed description. Note also that each dataset below is somewhat representative of one of our tasks of interest (with our synthetic data bucket being used for both conversational and background entity resolution).
@@ 4.1 Conversational Data
In this case, data is collected for entities that are relevant to the user interaction with the agent. To do this, annotators are shown sample conversations between a user and an agent with synthetic lists of entities provided, and asked to provide queries that unambiguously reference an arbitrarily picked entity in the aforementioned synthetic list. Annotators might thus be provided with a synthesized list of businesses or alarms and asked to refer to a particular entity within that list.
For example, the annotator might be shown a list of businesses that are synthetically constructed, and then asked to refer to a specific one in the list provided; for instance, they might say “Take me to the one that’s second from the bottom” or “Call the one on Main Street”.
@@ 4.2 Synthetic Data
The pipeline used to generate the synthetic data comprises of two parts: a set of templates and a list accompanying each template. The first part, a “language template”, contains different variations of queries that can be used for targeted cases, with slots present that can be filled pragmatically from those defined in the “slot list”. The second, a “slot list” accompanying the aforementioned template, includes mentions and other possible slot values (often comprising of named entities that aren’t mentions, or other slots that can take a large number of possible values such as date-times) if necessary. The slot list also contains the ground truth entity (or entities) that the mentions listed, when filled into the language template, could resolve to.
The data generation pipeline then takes the language template and slot list, and uses them to generate the possible queries. It does this by substituting corresponding values from the slot lists into the language templates to obtain fully formed user queries. The corresponding synthetic data is formed by using these queries and the ground truth entities present in the slot list, and adding in entities of other types into the data to serve as random negatives.
For example, a given language template might consist of phrases like “share [mention] with [name]” and “send [mention] to [name] please”. The corresponding slot list might have “[mention]” mapping to “this address” and “that address”, “[name]” mapping to various person names, and the ground truth entity tagged as “email address” and “physical address”. The pipeline then generates queries like “share that address with Mom” with “email address” and “physical address” entities marked as possible ground truth entity types, and entities of other types marked as negative.
@@
Figure 1: Sample screenshots used in the annotation of on-screen data. The data was annotated in a two-step process, as described in Section 4.3.
4.3 On-screen Data
As in Bhargava et al. (2023), screen data were collected from various web pages where phone number, e-mail and/or physical address information exist. Our on-screen data annotation comprised of a two-phase process. The first phase was used to obtain queries based on the screens shown, and the second one was for identifying the entities and mention for the given query. In the first grading project, annotators were given a screenshot (Figure 1(a)) with green and red boxes, and were asked to classify the green boxed data into one of the entities such as phone number, email address, etc. Then, annotators were then asked to provide three unique queries for the green boxed data.
In the second annotation project (Figure 1(b)), queries collected in the first step were shown to annotators one by one with their corresponding screenshots (but this time, without the bounding boxes), and with all the screen entities as a list. The annotators were asked if the query contains a mention to one of the given visual entities, and if the query sound natural. They were also asked to provide the entities from the list that were referred to in the given query, and to tag the part of the query referring that entity.
@@ 5 Models
We compare our proposed model ReALM, described in detail in Section 5.3 below, with two baseline approaches: one based on the reference resolver proposed in MARRS (Section 5.1), which is non-LLM based, and one based on ChatGPT (both GPT-3.5 and GPT-4; Section 5.2).
(a) Conversational User Turns
(b) Onscreen Capture
Figure 2: Technical diagrams representing user turns with a conversational assistant in (a), and a user screen in (b). Shaded rectangles represent various elements shown on the screen detectable by screen parser-extractors.
@@ 5.3 Our Approach
In this section, we provide examples of conversational and onscreen reference resolution tasks, followed by how we prompt the model to resolve the same.
@@
Select which among the following entities, if any, are required to understand the user request below. Output 0 if none of the entities are relevant. User request: Call the one on Rainbow St User Entities: 0. None 1. Type: Local Business | Name: Walgreens | Address: 225 Rainbow St, San Jose CA 94088 2. Type: Local Business | Name: CVS | Address: 105 E El Camino Real, Sunnyvale, CA 94087 3. Type: Local Business | Name: Qwark | Address: 1287 Hammerwood Ave, Sunnyvale, CA Relevant entity:
Select which among the following entities, if any, are required to understand the user request below. Output 0 if none of the entities are relevant. User request: Save the phone number at the bottom-right Screen: Your New home! Steven Realtors Inc. Trusted by over 5 million Proud homeowners Contact Us Monday -                          Saturday - Friday                              Sunday {{1. (206) 198 1999}}        {{2. (206) 198 1699}} Relevant entity:
Each data point consisting of a user query and the corresponding entities is converted into a sentence-wise format that we can feed to an LLM for training. Examples of the input before and after processing are shown in Appendix Sections A and C, with examples of how we convert entities of different types into text shown in Appendix B. Note that the entities are shuffled before being sent to the model so that the model does not overfit to particular entity positions.
With respect to the output that the model predicts, empirically, we find that the model is consistently able to predict a valid integer (or list of integers), without deviating and outputting any other text. In addition, we observe that the model also respects general output constraints (such as not predicting a ‘0‘ that represents ‘None of These‘ at the same time as one or more other entities) as well as those constrains enforced by the input (such as ensuring all predicted entity indices actually exist on the input side). The one exception that we observe is that, on occasion, we find that the model predicts the same entity twice (successively) in its output list. The only post-processing heuristic we apply is thus to convert the model’s predictions into a set of unique entities.
@@ 5.3.1 Conversational References
For the sake of this work, we assume conversational references to be of two types: type-based and descriptive. Type-based references are heavily reliant on using the user query in conjunction with the types of the entities to identify which entity (of a set of entities) are most relevant to the user query in question: for example, if the user says “play this”, we know that they are referring to an entity like a song or a movie, as opposed to a phone number or an address; “call him” likewise refers to a contact or possibly a phone number, as opposed to an alarm. Descriptive references, in contrast, tend to use a property of the entity to uniquely identify it: “The one in Times Square” for example might help uniquely refer to one among a set of addresses or business. Note that it is often the case that references might rely on both types and descriptions to unambiguously refer to a single object: consider the examples “play the one from Abbey Road” vs “directions to the one on Abbey Road”, both of which rely on both the entity type and description to identify a song in the first case and address in the second. In our proposed approach, we simply encode the type and various properties of the entity. We show our detailed encoding scheme in Appendix B.
@@ 5.3.2 Onscreen References
Data: List of turn objects
Result: Onscreen parse
1 $onscreen\_parse\leftarrow$ Empty list of onscreen parse elements;
2
// Step 0: Get all text boxes present in the screen
3 for each turn object $t$, index $i$ do
       // Step 1: Get unique surrounding objects
4       $surrounding\_objects\leftarrow$ Set of surrounding objects for $t$;
5
      // Step 2: Insert turn objects into the set
7
8
// Step 3: Sorting the centers of all surrounding objects
10
// Step 4: Determine vertical levels
11 $margin\leftarrow$ Margin for considering objects at the same level;
12 $levels\leftarrow$ List of vertical levels;
13 for each object $o$ in $sorted\_objects$ do
14       $same\_level\leftarrow$ List of objects at the same level as $o$;
15       for each object $other$ in $sorted\_objects$ do
17                   $same\_level\leftarrow same\_level\cup\{other\}$;
18
19
20      $levels\leftarrow levels\cup\{same\_level\}$;
21
22
// Step 5: Construct onscreen parse
23 for each level $l$ in $levels$ do
24       $level\_parse\leftarrow$ Empty string;
25       for each object $obj$ in $l$ do
[…]
27
[…]
29
30return $onscreen\_parse$;
Algorithm 1 Onscreen Parse Construction with Turn Object Injection
@@ 6 Results
Table 3: Model Accuracy for Different Datasets. A prediction is correct if the model correctly predicts all relevant entities, and incorrect otherwise. Conv refers to the Conversational Dataset, Synth to the Synthetic one, Screen to the Onscreen one and Unseen to a conversational dataset pertaining to a held-out domain.
@@
(a) Semantic Understanding
User Request: Call the evening Number
Screen:
{{1. 9 AM - 5 PM}}
{{2. 901.969.3120}}
{{3. 5 PM - 9 PM}}
{{4. 901.969.3391}}
Model Output: 4
(b) Summarisation
User Request: Remind me to get printouts
before the tax deadline
Screen:
Tax Deadlines 2023
{{1. Feb 15}}
Reclaim your tax exemption from withholding
{{2. April 18}}
First-quarter estimated tax payment due
Model Output: 2
(c) World Understanding
User Request: Take me to the one in Washington
Screen:
Indian Embassy
{{1. 1701 El Camino Real, Mountain View 94040}}
{{2. 333 Dexter Ave N, Seattle 98109}}
{{3. 8295 Tournament Drive, Memphis, TN 38125}}
Model Output: 2
(d) Commonsense Reasoning
User Request: Save the link to the breakfast Recipe
Screen:
IMAGE
Strawberry Granola
{{1. Recipe link}}
IMAGE
Lavender boba tea
{{2. Recipe link}}
Model Output: 1
Table 4: Qualitative examples that demonstrate the ability of ReALM to adapt to complex use-cases.
@@
Additionally, we also experiment with models of different sizes. We see that while performance in general improves across all dataset families with an increase in model size, the difference is most pronounced for the onscreen datasets, which alludes to the task being more complex in nature. Interestingly, and contrary to an otherwise consistent trend of larger models performing better, we find that performance on our Unseen dataset, which contains a held-out domain, first decreases with an increase in model size before increasing again. We hypothesize that this is due to the double-descent phenomenon Nakkiran et al. (2019).